OpenStudy (anonymous):

how to find the quadratic equation that has the vertex (-2,5) and whose graph passes through the point (0,9)

OpenStudy (anonymous):

\[f(x)=a(x-h)^{2}+k\] you know this formula right?

OpenStudy (anonymous):

haha nope slept through it... haha yea...

OpenStudy (anonymous):

read you class notes!! but plug in your vertex so you get \[f(x)=a(x+2)^{2}+5\] so far so good?

OpenStudy (anonymous):

haha dude he talk most the time and his notes are all over the place, and how i take notes while sleeping??

OpenStudy (anonymous):

oh and i had a game on that day too XD sooooo yea..

OpenStudy (anonymous):

but still did you understand?

OpenStudy (anonymous):

yea i get the vertex but how you know if it pass through the lines that was given??

OpenStudy (anonymous):

then yu plug in the point. \[9=a(0+2)^{2}+5\]

OpenStudy (anonymous):

oh okay. then you solve??? but solve for what?

OpenStudy (anonymous):

solve for a

OpenStudy (anonymous):

okay i got one and what you do with the answer??

OpenStudy (anonymous):

plug in yur a,h, and k values.

OpenStudy (anonymous):

ummmmm what??? it'll just be the same rifht?? f(x)=(x+2)^3+5?

OpenStudy (anonymous):

*right need my auto correct haha

OpenStudy (anonymous):

nope. its not a 3. its a 2 but yeah thats pretty much it.

OpenStudy (anonymous):

sorry that was bad typing skills...

OpenStudy (anonymous):

haha noticed.

OpenStudy (anonymous):

im doing chem while doing math... so yea... and my mom watching corny asian drama and the flirting hella stupid...

OpenStudy (anonymous):

haha how nice... did yu understand the answer tho?

OpenStudy (anonymous):

haha okay, and yes i got it. haha im super sleepy... need to finish homework... haha you start on chem dont forget

OpenStudy (anonymous):

haha finish yurs. ima start in the morning.

OpenStudy (anonymous):

haha yea i gotta catch up on apush im still on chap 8 TT.TT sad...

OpenStudy (anonymous):

haha then? why yu here? haha finish! haha jkjk